express the solution in exact form
ln(-x)+ ln 3=ln(3x-9)

Answer by stanbon(48535)

ln(-x)+ ln 3=ln(3x-9)
ln(-3x) = ln(3x-9
Take the anti-ln of both sides to get:
-3x = 3x-9
6x=9
x=3/2
---------
But ln(-3/2) is meaningless.
Therefore: No solutions.
==========
Cheers,
Stan H.